OpenStudy (anonymous):

A cube of gold has a side length of .050 cm. The mass of the cube is 2.40 kg. What is the density of the gold in g/cm3? Use significant figures

OpenStudy (anonymous):

density = mass/volume mass = 2.40 kg = 2.40 * 1000 g = 2400 g volume = volume of cube = (side length)^3 = (0.50 cm)^3 = 0.125 cm^3 2400 g density = ----------- = 19200 g/cm^3 0.125 cm^3